# A Survey on Diffusion Models for Time Series and Spatio-Temporal Data

### v3

@@ 4.2 Improved/Advanced Diffusion Model
We have already discussed the improvement methods of diffusion model in Sec. 2.3.3 and Sec. 2.3.4. In the following, we will analyze the methods from the more frequently used ones.
4.2.1 Conditional Diffusion Model
@@
Compared to standard approaches, conditional diffusion models are able to use given conditional information (e.g., different representations, different modalities, etc.) to steer the generative process and produce high-quality outputs that are tightly correlated with conditions [34]. More condition-based models are used for high-quality generation tasks.
@@
CSDI and MIDM use conditional diffusion models for time series imputation [56, 262]. They utilize score-based and probability-based diffusion models conditioned on observed values and through explicit imputation training. They can leverage the correlations among observed values to further enhance performance. Similar to CSDI, DiffAD and ImDiffusion employ a similar method based on conditional weight-incremental diffusion to enhance the imputation performance of missing values and are applied for time series anomaly detection [260, 257]. These approaches preserve the information of observed values and significantly improve the generation quality for stable anomaly detection. On the other hand, there are models that utilize different guiding information [266]. For example, [169] adjusts the diffusion model based on statistical information such as mean, standard deviation, Z-scores, and skewness, thereby synthesizing sensor data. [177] proposes two different conditional enhancement techniques to enable the model to adaptively consider conditional information a priori, thus performing the speech enhancement task. MEDiC [110] introduces a class-conditional DDPM approach to generate synthetic EEG embeddings. VGCDM [263] employs a pulse voltage-guided diffusion model along with a cross-attention mechanism for more efficient generation of electrical signals. Meanwhile, Wang et al. leverage multimodal information (images and text) as generative conditions to enhance the quality of real-time sales forecasts [264]. DiffShape [277] introduces a self-supervised diffusion learning mechanism that uses real sub-sequences as conditions. By leveraging a large amount of unlabeled data, it enhances the similarity between learned shapelets and actual sub-sequences for classification.
@@
For spatio-temporal data, most models adopt graph structures for representation, but the method of using conditions is similar to that of time series data [85, 184]. PriSTI draws inspiration from CSDI of conditioning on the observed value, employing a conditional feature extraction module based on linear interpolation for generating missing data [111]. DiffTraj estimates noise levels accurately using various external factors, such as the region of the trip and departure time [81]. ControlTraj further extends DiffTraj with the constraint of road network structures [290]. Additionally, many works guide their models using graph structures or feature graphs as conditions. For instance, DiffSTG [43] and USTD [170] use historical graph signals and graph structures as conditions, combining the spatio-temporal learning capabilities of GNNs and the uncertainty measurement of diffusion models to enhance the performance of forecasting tasks. DiffUFlow [35] utilizes extracted spatio-temporal feature graphs overlaid on coarse-grained flow graphs as a condition to guide the reverse process.
4.2.2 Latent Diffusion Model
@@
LDCast utilizes LDM for near-term precipitation forecasting, enhancing training stability, optimizing computational demands, and accurately representing the uncertainty of forecasting compared to GAN [93]. CLDM presents a conditional latent diffusion model based on latent space mapping, which breaks down the generative task into deterministic forecasting and the generation of predictive error scenarios, thus increasing efficiency within the latent space [107]. Similarly, Aristimunha et al. suggest the use of LDM for generating sleep EEG signals, where the latent feature maps outputted by encoders serve as input to the diffusion model, with the generated results obtained through the corresponding decoder [128]. In a similar manner, LADM proposes a method for spatio-temporal trajectory forecasting, where VAE acts as a generator and DDPM as a refiner [265]. Furthermore, Feng et al. develop a latent diffusion transformer for time series forecasting, aimed at compressing multivariate time stamp patterns into succinct latent representations and efficiently generating authentic multivariate time stamp values in a continuous latent space [253].
4.2.3 Other Variants of Diffusion Models
@@
There are also other methods that incorporate the different models above or use other more advanced diffusion models. For example, to accelerate generation, MedDiff for the first time utilizes DDIM with a new sampling strategy to generate high-dimensional large-scale electronic medical records [48]. Specifically, it incorporates a classifier to guide the sampling process, assigning high probabilities to the data with the correct labels. DCM is a diffusion-based causal inference model accelerated by DDIM, that more accurately captures counterfactual distributions under unmeasured confounding factors [267]. Asperti et al. propose GED, which uses conditional DDIM integrated with additional information and a post-processing network for high-performance weather forecasting [268].
@@
Besides, some algorithms have also been proposed that use a two-stage or a diffusion model to fuse with other models. TSDM adopts an improved two-stage diffusion model for identifying and reconstructing measurements with various uncertainties for power system measurement recovery, where the first stage includes a classifier-guided conditional anomaly detection component, and the second stage involves a diffusion-based measurement imputation component [269]. Mueller proposes using attention-enhanced condition-guided DDIM to tackle sample imbalance and data scarcity issues, along with simple DDPM for signal denoising, thereby synthesizing effective machine fault data for efficient anomaly detection [270]. Similarly, Wong et al. utilize DDIM for load domain adaptation, further enhanced with CNN for bearing fault anomaly detection [272].
5 Task Perspective
@@ 5.1 Forecasting
The field of time series forecasting has seen significant advancements with the incorporation of diffusion models. TimeGrad [13] and D3VAE [86] both employ diffusion probabilistic models to enhance forecasting, with TimeGrad focusing on autoregressive techniques (i.e., RNN) for probabilistic forecasting and D3VAE introducing a bidirectional variational auto-encoder that includes diffusion and denoising processes. This generative approach is further extended in the work by D-Va [47], which addresses the stochastic nature of stock price data through a deep hierarchical VAE combined with diffusion probabilistic techniques. Meanwhile, the study presented in [104] takes a different approach by modeling temporal data as continuous functions, allowing for the handling of irregularly sampled data. Building upon the diffusion model, TimeDiff [34] introduces novel conditioning mechanisms, future mixup, and autoregressive initialization, to improve time series forecasting. Lastly, the research in [164] explores task-agnostic unconditional diffusion models, proposing TSDiff, which employs a self-guidance mechanism for versatile time series applications.
@@
For spatio-temporal data, various diffusion model-based approaches have been proposed to tackle complex forecasting problems. DiffSTG [43] is the first work that generalizes denoising diffusion probabilistic models to spatio-temporal graphs, aiming to model complex spatio-temporal dependencies and intrinsic uncertainties within spatio-temporal graph data for better forecasting. DYffusion [285] introduces a framework that trains a stochastic, time-conditioned interpolator and a forecaster network to perform multi-step and long-range probabilistic forecasting for spatio-temporal data. On the other hand, DiffUFlow [35] focuses on urban data, which aims to address the challenge of fine-grained flow inference. DSTPP [70] provides a novel parameterization for spatio-temporal point processes. Meanwhile, SwinRDM [16] and DOT [65] demonstrate the adaptability of diffusion models in improving the quality of weather forecasts and travel time estimations, showcasing the wide applicability of these models across different domains and tasks. These works highlight the growing impact and potential of diffusion models in advancing spatio-temporal data analysis.
5.2 Generation
@@
Inspired by the powerful ability of the diffusion model in high-dimensional distribution learning, an intuitive usage is applying the learned diffusion model for data generation. Currently, a variety of diffusion models have been proposed for generating audio data. WaveGrad [20] is a pioneering work that utilizes gradient-based sampling to generate high-fidelity audio waveforms, marking a significant advancement in audio synthesis. Similarly, DiffWave [91] employs a non-autoregressive approach to achieve efficient and high-quality raw audio synthesis through a Markov chain process. Both WaveGrad and DiffWave are part of the same research lineage, leveraging the power of diffusion models to create complex waveforms from simple noise distributions. DOSE [177] takes a different approach by focusing on the speech enhancement task, which introduces a model-agnostic method that integrates condition information into the diffusion process.
Besides, diffusion model is also used in sequential recommendation. DiffuASR [23] proposes a diffusion-based sequence generation framework to address the data sparsity and long-tail user problems in sequential recommendation systems. It introduces a sequential U-Net designed for discrete sequence generation tasks and utilizes two guide strategies to assimilate preferences between generated and original sequences. DreamRec [167] employs a Transformer encoder to create guidance representations as the condition in the diffusion process.
@@
In the realm of spatial-temporal data, diffusion model is adopted to generate trajectory data. For instance, DiffTraj [81, 290] proposes the first effort that generates high-quality human trajectory with an unconditioned model, with the motivation to protect privacy. Meanwhile, [65] generates trajectories in the format of grids, by a condition model guided by the origin-destination information.
5.3 Imputation
@@
In the domain of time series and spatial-temporal data analysis, imputation refers to generating the unobserved data conditioned on the given observed data. CSDI [56] proposes a score-based diffusion model that probabilistically imputes missing time series and spatio-temporal data. MIDM [262] redefines the evidence lower bound (ELBO) for conditional diffusion models tailored for multivariate time series imputation, which ensures the consistency of observed and missing values. PriSTI [111] introduces a conditional diffusion framework specifically designed for spatio-temporal data imputation, using global context priors and geographic relationships to tackle scenarios with high missing data rates due to sensor failures. More recently, TabCSDI [33] and MissDiff [113] leverage diffusion models for imputing missing values in tabular data, each addressing distinct aspects of this complex challenge. TabCSDI manages categorical and numerical data effectively, providing a tailored solution to diverse data types common in tabular datasets. On the other hand, MissDiff enhances the diffusion model’s training process by introducing a novel masking technique for regression loss, which ensures consistent learning of data distributions and robustness against different missing data scenarios.
5.4 Anomaly Detection
@@
Anomaly detection aims to identify anomalies in the given time series or spatial-temporal data, which is quite practical and crucial for many real-world applications. A majority of works have been proposed for general time series anomaly detection. Initially, DiffAD [260] and ImDiffusion [257] both explored the synergy of imputation techniques with diffusion models for time series anomaly detection, enhancing the robustness of anomaly detection processes by accurately modeling complex dependencies. Concurrently, [259] and [258] use similar diffusion-based methodologies to address anomaly detection but apply different enhancements to improve performance and computational efficiency. Meanwhile, [261] and DDMT [99] tackled the specific challenges of instability and noise, implementing advanced diffusion reconstruction techniques to maintain accuracy in dynamic environments.
@@
Besides the above methods designed for general time series anomaly detection. There are other works that apply advanced diffusion techniques across various domains. For instance, Maat [37] anticipates performance metric anomalies in cloud services using a conditional denoising diffusion model to forecast metrics and detect anomalies. [270] enhances data synthesis for machine fault diagnosis using an attention mechanism within a conditional diffusion model framework. Diffusion-UDA [271] proposes a diffusion-based method for unsupervised domain adaptation in submersible fault diagnosis, leveraging diffusion processes to adapt domains for effective fault recognition.
6 Data Perspective

### v5

@@ 4.2. Improved/Advanced Diffusion Model
We have already discussed the improvement methods of the diffusion model in Sec. 2.3.3 and Sec. 2.3.4. In the following, we will analyze the methods of the more frequently used ones.
4.2.1. Conditional Diffusion Model
@@
Compared to standard approaches, conditional diffusion models are able to use given conditional information (e.g., different representations, different modalities, etc.) to steer the generative process and produce high-quality outputs that are tightly correlated with conditions (Shen and Kwok, 2023). More condition-based models are used for high-quality generation tasks.
@@
CSDI and MIDM use conditional diffusion models for time series imputation (Tashiro et al., 2021; Wang et al., 2023d). They utilize score-based and probability-based diffusion models conditioned on observed values and through explicit imputation training. They can leverage the correlations among observed values to further enhance the performance. Similar to CSDI, DiffAD and ImDiffusion employ a similar method based on conditional weight-incremental diffusion to enhance the imputation performance of missing values and are applied for time series anomaly detection (Xiao et al., 2023; Chen et al., 2023c). These approaches preserve the information on the observed values and significantly improve the generation quality for stable anomaly detection. On the other hand, there are models that utilize different guiding information (Fu et al., 2024). For example, (Zuo et al., 2023) adjusts the diffusion model based on statistical information such as mean, standard deviation, Z-scores, and skewness, thereby synthesizing sensor data. (Tai et al., 2024) proposes two different conditional enhancement techniques to enable the model to adaptively consider conditional information a priori, thus performing the speech enhancement task. MEDiC (Sharma et al., 2023) introduces a class-conditional DDPM approach to generate synthetic EEG embeddings. VGCDM (Liu et al., 2023a) employs a pulse voltage-guided diffusion model along with a cross-attention mechanism for more efficient generation of electrical signals. Meanwhile, Wang et al. leverage multimodal information (images and text) as generative conditions to enhance the quality of real-time sales forecasts (Wang, 2024). DiffShape (Liu et al., 2024b) introduces a self-supervised diffusion learning mechanism that uses real sub-sequences as conditions. By leveraging a large amount of unlabeled data, it enhances the similarity between learned shapelets and actual sub-sequences for classification.
@@
For spatio-temporal data, most models adopt graph structures for representation, but the method of using conditions is similar to that of time series data (Ling et al., 2024; Liu et al., 2024a). PriSTI draws inspiration from CSDI of conditioning on the observed value, employing a conditional feature extraction module based on linear interpolation for generating missing data (Liu et al., 2023c). DiffTraj estimates noise levels accurately using various external factors, such as the region of the trip and departure time (Zhu et al., 2024a). ControlTraj further extends DiffTraj with the constraint of road network structures (Zhu et al., 2024b). Additionally, many works guide their models using graph structures or feature graphs as conditions. For instance, DiffSTG (Wen et al., 2023) and USTD (Hu et al., 2023) use historical graph signals and graph structures as conditions, combining the spatio-temporal learning capabilities of GNNs and the uncertainty measurement of diffusion models to enhance the performance of forecasting tasks. DiffUFlow (Zheng et al., 2023) utilizes extracted spatio-temporal feature graphs overlaid on coarse-grained flow graphs as a condition to guide the reverse process.
4.2.2. Latent Diffusion Model
@@
LDCast utilizes LDM for near-term precipitation forecasting, enhancing training stability, optimizing computational demands, and accurately representing the uncertainty of forecasting compared to GAN (Leinonen et al., 2023). CLDM presents a conditional latent diffusion model based on latent space mapping, which breaks down the generative task into deterministic forecasting and the generation of predictive error scenarios, thus increasing efficiency within the latent space (Dong et al., 2023). Similarly, Aristimunha et al. suggest the use of LDM for generating sleep EEG signals, where the latent feature maps outputted by the encoders serve as input to the diffusion model, with the generated results obtained through the corresponding decoder (Aristimunha et al., 2023). In a similar manner, LADM proposes a method for spatio-temporal trajectory forecasting, where VAE acts as a generator and DDPM as a refiner (Lv et al., 2024). Furthermore, Feng et al. develop a latent diffusion transformer for time series forecasting, aimed at compressing multivariate time stamp patterns into succinct latent representations and efficiently generating authentic multivariate time stamp values in a continuous latent space (Feng et al., 2024).
4.2.3. Other Variants of Diffusion Models
@@
There are also other methods that incorporate the different models above or use other more advanced diffusion models. For example, to accelerate generation, MedDiff for the first time utilizes DDIM with a new sampling strategy to generate high-dimensional large-scale electronic medical records (He et al., 2023). Specifically, it incorporates a classifier to guide the sampling process, assigning high probabilities to the data with the correct labels. DCM is a diffusion-based causal inference model accelerated by DDIM, that more accurately captures counterfactual distributions under unmeasured confounding factors (Shimizu, 2023). Asperti et al. propose GED, which uses conditional DDIM integrated with additional information and a post-processing network for high-performance weather forecasting (Asperti et al., 2023).
@@
Besides, some algorithms have also been proposed that use a two-stage or a diffusion model to fuse with other models. TSDM adopts an improved two-stage diffusion model for identifying and reconstructing measurements with various uncertainties for power system measurement recovery, where the first stage includes a classifier-guided conditional anomaly detection component, and the second stage involves a diffusion-based measurement imputation component (Pei et al., 2023). Mueller proposes using attention-enhanced condition-guided DDIM to tackle sample imbalance and data scarcity issues, along with simple DDPM for signal denoising, thereby synthesizing effective machine fault data for efficient anomaly detection (Mueller, 2024). Similarly, Wong et al. utilize DDIM for load domain adaptation, further enhanced with CNN for bearing fault anomaly detection (Wong et al., 2024).
4.3. Denoisers in the Diffusion Model
Denoisers in the diffusion model serve as the core modules that transform noise into structured signals, with their architectures, inductive biases, and prior-guided designs fundamentally shaping the model’s generative capacity, temporal reasoning, and applicability to diverse inverse problems. In the following, we will provide a detailed analysis of these aspects respectively.
4.3.1. Architecture of the Denoiser
The denoiser serves as the backbone of diffusion models, determining their expressive power and cross-modal generalization. U-Net remains the architecture most widely used in diffusion models (Ronneberger et al., 2015; Ho et al., 2020; Rombach et al., 2022), where its encoder–decoder design with skip connections supports multi-scale representation learning and effective propagation of global and local information. Models such as DiffuASR (Liu et al., 2023g) and DiffAD (Xiao et al., 2023) adapt U-Net with 1D convolutions or temporal dilations to better model sequential dependencies. Some works explore attention-based backbones for long-range dependency modeling, including CSDI (Tashiro et al., 2021) and PriSTI (Liu et al., 2023c), which enhance global context capture and scalability (Zheng et al., 2023). Hybrid architectures further extend denoising to structured data. D3R combines attention and diffusion for anomaly detection (Wang et al., 2024c), while SpecSTG uses graph spectral convolutions for traffic forecasting (Lin et al., 2024). These domain-aware designs show that integrating prior knowledge improves denoising efficiency and interpretability (Wen et al., 2023). Recent trends favor modular, cross-modal and lightweight denoisers, including foundation backbones with temporal or structural adaptations (Ma et al., 2025).
4.3.2. Inductive Bias in Temporal Denoisers
The effectiveness of diffusion models relies strongly on the inductive biases in their denoiser architectures. U-Net structures leverage translation equivariance and hierarchical feature extraction for robust spatial representations (Kadkhodaie et al., 2024; Kamb and Ganguli, 2025). For time series, analogous inductive priors must capture temporal continuity, periodicity, and causal dynamics. Hierarchical temporal architectures (e.g., Temporal U-Net, WaveGrad (Chen et al., 2020)) address multi-scale temporal patterns through progressive downsampling and upsampling. In spatio-temporal settings, CNN-based denoisers (e.g., DiffTraj (Zhu et al., 2024a)) learn long-term dependencies efficiently with skip connections, while graph-based denoisers (e.g., DiffSTG (Wen et al., 2023), DiffUFlow (Zheng et al., 2023)) incorporate structural priors via graph convolutions. Emerging temporal transformers (e.g., DiffUFlow (Zheng et al., 2023), PriSTI (Liu et al., 2023c)) unify causal and hierarchical biases through self-attention with relative positional encodings.
4.3.3. Denoiser-as-Prior in Temporal Inverse Problems
Beyond the full diffusion sampling framework, many inverse problems can be solved by directly leveraging the denoiser (or score function) as a learned prior. This idea has been widely explored under the Plug-and-Play and Score Prior paradigms, where the denoiser acts as a proximal operator for iterative reconstruction (Kadkhodaie and Simoncelli, 2020; Milanfar and Delbracio, 2025). For the time series field, a similar trend has emerged. Models such as ImDiffusion (Chen et al., 2023c) and SADI (Islam et al., 2025) employ the denoiser alone to restore missing segments, achieving high-quality reconstructions without full stochastic sampling. This approach reduces computational overhead and enhances interpretability by decoupling temporal prior learning from generative sampling. The denoiser-as-prior framework therefore provides a promising bridge between generative diffusion modeling and classical temporal signal restoration.
5. Task Perspective
@@ 5.1. Forecasting
The field of time series forecasting has seen significant advancements with the incorporation of diffusion models. TimeGrad (Rasul et al., 2021) and D3VAE (Li et al., 2022a) both employ diffusion probabilistic models to enhance forecasting, with TimeGrad focusing on autoregressive techniques (i.e., RNN) for probabilistic forecasting and D3VAE introducing a bidirectional variational auto-encoder that includes diffusion and denoising processes. This generative approach is further extended in the work by D-Va (Koa et al., 2023), which addresses the stochastic nature of stock price data through a deep hierarchical VAE combined with diffusion probabilistic techniques. Meanwhile, the study presented in (Biloš et al., 2022) takes a different approach by modeling temporal data as continuous functions, allowing the handling of irregularly sampled data. Building upon the diffusion model, TimeDiff (Shen and Kwok, 2023) introduces novel conditioning mechanisms, future mixup, and autoregressive initialization, to improve time series forecasting. Lastly, the research in (Kollovieh et al., 2024) explores task-agnostic unconditional diffusion models, proposing TSDiff, which employs a self-guidance mechanism for versatile time series applications.
@@
For spatio-temporal data, various diffusion model-based approaches have been proposed to tackle complex forecasting problems. DiffSTG (Wen et al., 2023) is the first work that generalizes denoising diffusion probabilistic models to spatio-temporal graphs, aiming to model complex spatio-temporal dependencies and intrinsic uncertainties within spatio-temporal graph data for better forecasting. DYffusion (Rühling Cachay et al., 2024) introduces a framework that trains a stochastic, time-conditioned interpolator and a forecaster network to perform multi-step and long-range probabilistic forecasting for spatio-temporal data. On the other hand, DiffUFlow (Zheng et al., 2023) focuses on urban data, which aims to address the challenge of fine-grained flow inference. DSTPP (Yuan et al., 2023a) provides a novel parameterization for spatio-temporal point processes. Meanwhile, SwinRDM (Chen et al., 2023a) and DOT (Lin et al., 2023b) demonstrate the adaptability of diffusion models in improving the quality of weather forecasts and travel time estimations, showcasing the wide applicability of these models across different domains and tasks. These works highlight the growing impact and potential of diffusion models in advancing spatio-temporal data analysis.
5.2. Generation
@@
Inspired by the powerful ability of the diffusion model in high-dimensional distribution learning, an intuitive usage is applying the learned diffusion model for data generation. Currently, a variety of diffusion models have been proposed for generating audio data. WaveGrad (Chen et al., 2020) is a pioneering work that utilizes gradient-based sampling to generate high-fidelity audio waveforms, marking a significant advancement in audio synthesis. Similarly, DiffWave (Kong et al., 2020) employs a non-autoregressive approach to achieve efficient and high-quality raw audio synthesis through a Markov chain process. Both WaveGrad and DiffWave are part of the same research lineage, leveraging the power of diffusion models to create complex waveforms from simple noise distributions. DOSE (Tai et al., 2024) takes a different approach by focusing on the speech enhancement task, which introduces a model-agnostic method that integrates condition information into the diffusion process. Besides, diffusion model is also used in sequential recommendation. DiffuASR (Liu et al., 2023g) proposes a diffusion-based sequence generation framework to address data sparsity and long-tail user problems in sequential recommendation systems. It introduces a sequential U-Net designed for discrete sequence generation tasks and utilizes two guide strategies to assimilate preferences between generated and original sequences. DreamRec (Yang et al., 2024b) employs a Transformer encoder to create guidance representations as conditions in the diffusion process.
@@
In the realm of spatio-temporal data, diffusion models are adopted to generate trajectory data. For instance, DiffTraj (Zhu et al., 2024a, b) proposes the first effort that generates high-quality human trajectories with an unconditioned model, with the motivation to protect privacy. Meanwhile, (Lin et al., 2023b) generates trajectories in the format of grids, by a condition model guided by the origin-destination information.
5.3. Imputation
@@
In the domain of time series and spatio-temporal data analysis, imputation refers to generating the unobserved data conditioned on the given observed data. CSDI (Tashiro et al., 2021) proposes a score-based diffusion model that probabilistically imputes missing time series and spatio-temporal data. MIDM (Wang et al., 2023d) redefines the evidence lower bound (ELBO) for conditional diffusion models tailored for multivariate time series imputation, which ensures the consistency of observed and missing values. PriSTI (Liu et al., 2023c) introduces a conditional diffusion framework specifically designed for spatio-temporal data imputation, using global context priors and geographic relationships to tackle scenarios with high missing data rates due to sensor failures. More recently, TabCSDI (Zheng and Charoenphakdee, 2022) and MissDiff (Ouyang et al., 2023) leverage diffusion models for imputing missing values in tabular data, each addressing distinct aspects of this complex challenge. TabCSDI manages categorical and numerical data effectively, providing a tailored solution to diverse data types common in tabular datasets. On the other hand, MissDiff enhances the diffusion model’s training process by introducing a novel masking technique for regression loss, which ensures consistent learning of data distributions and robustness against different missing data scenarios.
5.4. Anomaly Detection
@@
Anomaly detection aims to identify anomalies in the given time series or spatio-temporal data, which is quite practical and crucial for many real-world applications. A majority of works have been proposed for general time series anomaly detection. Initially, DiffAD (Xiao et al., 2023) and ImDiffusion (Chen et al., 2023c) both explored the synergy of imputation techniques with diffusion models for time series anomaly detection, enhancing the robustness of anomaly detection processes by accurately modeling complex dependencies. Concurrently, (Pintilie et al., 2023) and (Livernoche et al., 2023) use similar diffusion-based methodologies to address anomaly detection but apply different enhancements to improve performance and computational efficiency. Meanwhile, (Wang et al., 2024c) and DDMT (Yang et al., 2023b) tackled the specific challenges of instability and noise, implementing advanced diffusion reconstruction techniques to maintain accuracy in dynamic environments.
@@
Besides the above methods designed for general time series anomaly detection. There are other works that apply advanced diffusion techniques across various domains. For instance, Maat (Lee et al., 2023) anticipates performance metric anomalies in cloud services using a conditional denoising diffusion model to forecast metrics and detect anomalies. (Mueller, 2024) enhances data synthesis for machine fault diagnosis using an attention mechanism within a conditional diffusion model framework. Diffusion-UDA (Zhao et al., 2024c) proposes a diffusion-based method for unsupervised domain adaptation in submersible fault diagnosis, leveraging diffusion processes to adapt domains for effective fault recognition.
6. Data Perspective